# Mining Constraints from Reference Process Models for Detecting Best-Practice Violations in Event Logs

### v1

@@ 2 Motivation
$\sigma=\langle$create order, enter goods receipt, enter invoice, check invoice amount, make $\rangle$
@@ 3.2 Process Models and Process Model Collections
Using $\mathcal{M}$ to denote a process model collection, each process model $M\in\mathcal{M}$ is a tuple $M=$($A$, $F$, $R$, $D$). $A$ is the set of labeled steps in $M$, including tasks and events in a BPMN diagram or transitions in a Petri net. Just like for events, we call $a\in A\subseteq\mathcal{A}$ an activity in the remainder. $F$ is the set of unique finite execution sequences allowed by the model, with each $\pi\in F=\langle a_{1},...,a_{n}\rangle$ as a sequence of activities in $A$. Focusing on finite traces captures the intuition that each process execution is expected to complete in a finite number of steps [6]. $R\subseteq\mathcal{R}$ is a set of roles involved in the process represented by $M$ and, finally, $D:\mathcal{A}\rightarrow\mathcal{R}\cup\{\bot\}$ is a mapping that associates activities with roles, e.g., a create invoice activity may be mapped to a vendor role in a procurement process model.
@@ 3.3 Declarative Constraints
AlternatePrecedence(a,b) ($(\neg b$ S $a)\land$G $(b\rightarrow(\neg b$ S $a))$ b
@@
[…]
@@ 4.1.1 Coverage
Note that the constraints of all types we cover have a $support$ that indicates how often a constraint was extracted from the model collection $\mathcal{M}$ and is set when refining the collection of mined constraints $C_{\mathcal{M}}$.
@@
[…]
@@ 4.1.2 Constraint Extraction
Activity-level constraints. We extract activity-level constraints from a model $M=(A,F,R,D)$ by first instantiating potential constraints using each template from Table 3 and each pair-wise combination of activities in $A$. Then we check which of these potential constraints is satisfied by all and activated by at least one of the execution sequences in $F$. For instance, for the Response template, we check for each pair ($a_{1}$,$a_{2}$) $\in A\times A$ if G($a_{1}\rightarrow$ F $a_{2}$) (cf. Section 3) holds for all $\pi\in F$ and—if so—set $c_{\mathcal{M}}.type=activity$, $c_{\mathcal{M}}.templ=\textsc{Response}$, $c_{\mathcal{M}}.a_{1}=a_{1}$, and $c_{\mathcal{M}}.a_{2}=a_{2}$ and add $c_{\mathcal{M}}$ to $C_{M}$.
@@
[…]
@@
[…]
@@
t Output $res$: set containing a set of action sequences per business object
1:$res\leftarrow\{F_{o}\leftarrow\emptyset\mid o\in O\}$triangleright Initialize an empty set of action sequences for each business object
@@
8:                 $\pi_{o}\leftarrow\pi_{o}+\langle n_{a_{i}}\rangle$triangleright Retain the action applied to $o$ and append to current sequence
@@
10:return $res$
[…]
@@ 4.2.1 Constraint Fitting
Given a mined constraint $c_{\mathcal{M}}$ and an event log $L$, a fitted constraint $c_{L}$ is a tuple $c_{L}=(c_{\mathcal{M}},$ $type,$ $templ,$ $[components],$ $support,$ $sim,$ $relevance)$, with $c_{L}$’s $type$, $templ$, and $support$ being equal to those of $c_{\mathcal{M}}$. $c_{L}.[components]$ is a shorthand to refer to the list of $c_{L}$’s type-specific components (e.g., $c_{L}.[components]=c_{L}.a_{1},c_{L}.a_{2}$ for activity-level constraints) that are each linked to components (i.e., activities, actions, objects, or roles) found in $L$. Finally, $c_{L}.sim$ is a map capturing the semantic similarity between $c_{\mathcal{M}}$’s and $c_{L}$’s components, and $c_{L}.relevance$ is a score indicating the relevance of $c_{L}$ for $L$.
@@
Semantic similarity computation. A key part of Algorithm 2 is the match($x_{\mathcal{M}}$, $x_{L}$) function, which is used to decide whether a constraint component $x_{\mathcal{M}}$ is semantically similar to an event log component $x_{L}$. It is instantiated using a similarity measure sim and a threshold $\epsilon$, i.e., match($x_{\mathcal{M}}$, $x_{L}$)$:=\texttt{sim}(x_{\mathcal{M}},x_{L})>\epsilon$. As a similarity measure, we employ the cosine similarity based on embedding vectors for $x_{\mathcal{M}}$ and $x_{L}$, which we obtain from a pre-trained sentence transformer [11]. These transformers are specifically designed to capture semantically meaningful representations on the level of sentences rather than individual words and can thus be applied on activities or business objects that may consist of multiple words. We set $\epsilon$ to 0.5 by default, ensuring that all fitted constraints are reasonably similar to the components of $L$, which avoids having to filter out a lot of constraints from $C_{L}^{S}$ in the next step. [^3] Note that, in order to make match more or less restrictive, it is possible to use different values for $\epsilon$ depending on the type of component we aim to find matches for, i.e., setting different thresholds for activities, objects, and actions.
@@
3:    if $c_{\mathcal{M}}.type$ is $activity$ then
@@
7:    if $c_{\mathcal{M}}.{type}$ is $interobj$ then
@@
11:    if $c_{\mathcal{M}}.type$ is $intraobj$ then
@@
22:    if $c_{\mathcal{M}}.type$ is $role$ then
@@
For instance, consider a constraint $c_{\mathcal{M}}$ corresponding to AtLeastOne(check) $\mid$ order and an event log that relates to purchase order objects, including an examine purchase order activity. Then, our framework would recognize that order and purchase order are semantically similar, i.e., match(order, purchase order)$=true$, and that examine is a synonym of check, i.e., syn(check, examine)$=true$, yielding a fitted constraint $c_{L}$ corresponding to AtLeastOne(examine) $\mid$ purchase order.
@@ 4.3.1 Trace-based Conformance Checking
The goal of this step is to check which constraints are violated per trace $\sigma\in L$. To this end, our framework creates a collection $V_{L}$ of trace-constraint pairs ($\sigma$,$c_{L}$), where each pair represents the violation of $c_{L}$ in $\sigma$. To establish these pairs, our framework checks per trace $\sigma\in L$ and each constraint $c_{L}\in C_{L}^{S}$ if that constraint is violated, i.e., if $\sigma\not\models c_{L}$. The checking process is analogous to the one used in constraint mining (cf. Section 4.1). However, while in constraint mining we assess whether a constraint is fulfilled across all sequences allowed by a model, here we focus on identifying specific traces where a constraint is not met. Given a trace $\sigma$, the conformance checks per constraint type are performed as follows:
@@ 4.3.2 Result Representation
AtLeastOne(create)||||order Each order must be created $\{\sigma_{6},\dots,\sigma_{76}\}$
Precedence(check, approve)||||invoice An invoice must be checked before it is approved $\{\sigma_{23}\}$
@@ 5.1 Data Collection
Original logs 1,500 4.9 18 3.9 17 3.5 $720$ 3.7 10
Noisy logs 1,500 4.8 18 3.9 17 30.9 $703$ 3.9 14
@@ 5.3 Results
After refinement, i.e., after equal or highly similar constraints were standardized and weaker, redundant ones were filtered out, 55,913 activity-level, 20,920 inter-object, 8,223 intra-object constraints, and 2,387 role constraints remain, as depicted in Table 8. Many constraints were extracted from multiple models, as shown by the number of constraints with a support of $>1$. Notably, more than half of the intra-object constraints (4,484) were extracted from multiple models, which suggests that these are more generally applicable than the other types of constraints, for which this ratio is lower.
@@
Const. type Count Support $>1$

### v2

@@ 2 Motivation
[…]
@@ 3.2 Process Models and Process Model Collections
Using $\mathcal{M}$ to denote a process model collection, each process model $M\in\mathcal{M}$ is a tuple M=M=(AA, $F$, $R$, $D$). AA is the set of labeled steps in $M$, including tasks and events in a BPMN diagram or transitions in a Petri net. Just like for events, we call $a\in A\subseteq\mathcal{A}$ an activity in the remainder. $F$ is the set of unique finite execution sequences allowed by the model, with each $\pi\in F=\langle a_{1},...,a_{n}\rangle$ as a sequence of activities in $A$. Focusing on finite traces captures the intuition that each process execution is expected to complete in a finite number of steps [6]. $R\subseteq\mathcal{R}$ is a set of roles involved in the process represented by $M$ and, finally, $D:\mathcal{A}\rightarrow\mathcal{R}\cup\{\bot\}$ is a mapping that associates activities with roles, e.g., a create invoice activity may be mapped to a vendor role in a procurement process model.
@@ 3.3 Declarative Constraints
AlternatePrecedence(a,b) $(\neg b$ S $a)\landG$ $(b\rightarrow(\neg b$ S $a))$ b
@@
NotCoExistence(a,b) (F a $\rightarrow\negF$ b) $\land$ (F b $\rightarrow\negF$ a) a,b
@@ 4.1.1 Coverage
Note that the constraints of all types we cover have a supportsupport that indicates how often a constraint was extracted from the model collection $\mathcal{M}$ and is set when refining the collection of mined constraints $C_{\mathcal{M}}$.
@@
An intra-object constraint $c_{\mathcal{M}}\in C_{\mathcal{M}}$ is a tuple $c_{\mathcal{M}}=(type,templ,$ $object,arity,n_{1},n_{2},support)$, with $c_{\mathcal{M}}.type=intraobj$, $c_{\mathcal{M}}.templ$ the constraint’s Declare template, $c_{\mathcal{M}}.object$ its object, and $c_{\mathcal{M}}.arity\in\{\emph{unary},\emph{binary}\}$ its arity indicating whether $c_{\mathcal{M}}$ is a unary or binary constraint, according to $c_{\mathcal{M}}.templ$. Moreover, $c_{\mathcal{M}}.n_{1}$ corresponds to the constraint’s first action, $c_{\mathcal{M}}.n_{2}$ represents its second action (set to $\bot$ if $c_{\mathcal{M}}.arity=\text{unary}$), and $c_{\mathcal{M}}.support$ pertains to its support.
@@ 4.1.2 Constraint Extraction
Activity-level constraints. We extract activity-level constraints from a model $M=(A,F,R,D)$ by first instantiating potential constraints using each template from Table 3 and each pair-wise combination of activities in $A$. Then we check which of these potential constraints is satisfied by all and activated by at least one of the execution sequences in $F$. For instance, for the Response template, we check for each pair ($a_{1}$$a_{2}$) $\in A\times A$ if G($a_{1}\rightarrow$ F $a_{2}$) (cf. Section 3) holds for all $\pi\in F$ and—if so—set $c_{\mathcal{M}}.type=activity$, $c_{\mathcal{M}}.templ=\textsc{Response}$, $c_{\mathcal{M}}.a_{1}=a_{1}$, and $c_{\mathcal{M}}.a_{2}=a_{2}$ and add $c_{\mathcal{M}}$ to $C_{M}$.
@@
To obtain $F_{obj}$, we project each (activity) execution sequence $\pi\in F$ to an object sequence $\pi_{\emph{obj}}$, which we achieve using a semantic extraction technique [9] to obtain the business objects (if any) from each activity in $\pi$. For instance, for $\pi=\langle\emph{create order},\emph{check order},\emph{approve order},\emph{ship goods}\rangle$, we get $\pi_{\emph{obj}}=\langle\emph{order},\emph{order},\emph{order},\emph{goods}\rangle$. If an activity does not relate to any business object, it is not considered in the projected trace. While uncommon, in cases where an activity relates to multiple business objects, these are considered as a single business object, assuming that in a reference process model, these are then consistently referenced together.
Afterwards, we instantiate constraint templates between possible pairs of objects $(o_{1},o_{2})$ occurring in $F_{\emph{obj}}$ and check them against the sequences in $F_{\emph{obj}}$, in the same manner as done for activity-level constraints. For instance, for the binary RespondedExistence template, we check for each combination ($o_{1}$$o_{2}$) with $o_{1},o_{2}\in O$ if F $o_{1}$ $\rightarrow$ F $o_{2}$ holds for all $\pi_{\emph{obj}}\in F_{\emph{obj}}$ and—if so—set $c_{\mathcal{M}}.type=interobj$, $c_{\mathcal{M}}.templ=\textsc{RespondedExistence}$, $c_{\mathcal{M}}.o_{1}=o_{1}$, and $c_{\mathcal{M}}.o_{2}=o_{2}$ and add $c_{\mathcal{M}}$ to $C_{M}$.
@@
To obtain the necessary information on business objects and actions, we, therefore, again use the semantic extraction technique [9] (as done for inter-object constraints). This time, we turn the output of that approach into a function getObjectActionPairs, which, given an activity $a$, returns a set of object-type-action pairs obtained from $a$, where each pair is given as a tuple ($o$, $n$), indicating that action $n$ is applied to object oo. [^2] Note that, for clarity, we here focus on situations where each activity yields a single object-action pair, although our framework can also handle activities with multiple pairs, e.g., receive and check document. For instance, $\texttt{getObjectActionPairs}(\emph{approve purchase order})=\{(\emph{purchase order},\emph{approve})\}$. By applying the function to each activity $a\in A$, we obtain a set $O$ of distinct business objects that appear in the activities in $A$ and a set $N_{o}$ of actions per business object $o\in O$, which we use to create intra-object sequences.
As shown in Algorithm 1, we create a set of intra-object sequences $F_{o}$ for each business object $o\in O$ as follows: For each $\pi\in F$, we check for each activity $a_{i}$ in $\pi$ if $a_{i}$ refers to $o$ and if so only retain the action applied to $o$. For instance, for $\pi=\langle\emph{create order},\emph{check order},\emph{approve order},\emph{ship goods}\rangle$, we get $\pi_{\emph{order}}=\langle\emph{create},\emph{check},\emph{approve}\rangle$.
@@
t Output resres: set containing a set of action sequences per business object
[…]
@@
8:      $\pi_{o}\leftarrow\pi_{o}+\langle n_{a_{i}}\rangle$ $\triangleright$ Retain the action applied to $o$ and append to current sequence
@@
10:return resres
Based on an intra-object set of sequences $F_{o}$, we can then instantiate the Declare templates with actions in $N_{o}$ in the same manner as done with entire activities for activity-level constraints. For instance, for the binary template Precedence, we check for each combination $(n_{1},n_{2})\in N_{o}\times N_{o}$ if G($n_{2}\rightarrow$ O $n_{1}$)(cf. Section 3) holds for all $\pi_{o}\in F_{o}$ and—if so—set $c_{\mathcal{M}}.type=intraobj$, $c_{\mathcal{M}}.templ=\textsc{Precedence}$, $c_{\mathcal{M}}.arity=\text{binary}$, $c_{\mathcal{M}}.n_{1}=n_{1}$, $c_{\mathcal{M}}.n_{2}=n_{2}$, and $c_{\mathcal{M}}.object=\emph{order}$ and add $c_{\mathcal{M}}$ to $C_{M}$.
@@ 4.2.1 Constraint Fitting
Given a mined constraint $c_{\mathcal{M}}$ and an event log $L$, a fitted constraint $c_{L}$ is a tuple $c_{L}=(c_{\mathcal{M}},$ $type,$ $templ,$ $[components],$ $support,$ $sim,$ $relevance)$, with $c_{L}$’s typetype, templtempl, and supportsupport being equal to those of $c_{\mathcal{M}}$. $c_{L}.[components]$ is a shorthand to refer to the list of $c_{L}$’s type-specific components (e.g., $c_{L}.[components]=c_{L}.a_{1},c_{L}.a_{2}$ for activity-level constraints) that are each linked to components (i.e., activities, actions, objects, or roles) found in $L$. Finally, $c_{L}.sim$ is a map capturing the semantic similarity between $c_{\mathcal{M}}$’s and $c_{L}$’s components, and $c_{L}.relevance$ is a score indicating the relevance of $c_{L}$ for $L$.
@@
[…]
@@
3:  if $c_{\mathcal{M}}.type$ is activityactivity then
@@
7:  if $c_{\mathcal{M}}.{type}$ is interobjinterobj then
@@
11:  if $c_{\mathcal{M}}.type$ is intraobjintraobj then
@@
22:  if $c_{\mathcal{M}}.type$ is rolerole then
@@
For instance, consider a constraint $c_{\mathcal{M}}$ corresponding to AtLeastOne(check) $\mid$ order and an event log that relates to purchase order objects, including an examine purchase order activity. Then, our framework would recognize that order and purchase order are semantically similar, i.e., match(order, purchase order)=true=true, and that examine is a synonym of check, i.e., syn(check, examine)=true=true, yielding a fitted constraint $c_{L}$ corresponding to AtLeastOne(examine) $\mid$ purchase order.
@@ 4.3.1 Trace-based Conformance Checking
The goal of this step is to check which constraints are violated per trace $\sigma\in L$. To this end, our framework creates a collection $V_{L}$ of trace-constraint pairs ($\sigma$$c_{L}$), where each pair represents the violation of $c_{L}$ in $\sigma$. To establish these pairs, our framework checks per trace $\sigma\in L$ and each constraint $c_{L}\in C_{L}^{S}$ if that constraint is violated, i.e., if $\sigma\not\models c_{L}$. The checking process is analogous to the one used in constraint mining (cf. Section 4.1). However, while in constraint mining we assess whether a constraint is fulfilled across all sequences allowed by a model, here we focus on identifying specific traces where a constraint is not met. Given a trace $\sigma$, the conformance checks per constraint type are performed as follows:
@@ 4.3.2 Result Representation
AtLeastOne(create)||order Each order must be created $\{\sigma_{6},\dots,\sigma_{76}\}$
Precedence(check, approve)||invoice An invoice must be checked before it is approved $\{\sigma_{23}\}$
@@ 5.1 Data Collection
Original logs 1,500 4.9 18 3.9 17 3.5 720720 3.7 10
Noisy logs 1,500 4.8 18 3.9 17 30.9 703703 3.9 14
@@ 5.3 Results
After refinement, i.e., after equal or highly similar constraints were standardized and weaker, redundant ones were filtered out, 55,913 activity-level, 20,920 inter-object, 8,223 intra-object constraints, and 2,387 role constraints remain, as depicted in Table 8. Many constraints were extracted from multiple models, as shown by the number of constraints with a support of >1>1. Notably, more than half of the intra-object constraints (4,484) were extracted from multiple models, which suggests that these are more generally applicable than the other types of constraints, for which this ratio is lower.
@@
Const. type Count Support >1>1